OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality 5x + y greater than or equal to -10. PLEASE HELP !!!!!!!!!!

OpenStudy (harsimran_hs4):

our first step in such situations is to plot 5x + y = -10 line (which is the boundary line also) so can you draw that line and show it ?

OpenStudy (anonymous):

no .. i'm really confused @harsimran_hs4

OpenStudy (harsimran_hs4):

ok i`ll draw the line here`s it so can you tell if (0,0) satisfies this inequality?

OpenStudy (anonymous):

No it doesn't satisfies the inequality. so it would be a Solid line with shading on the side that includes the origin, correct? @harsimran_hs4

OpenStudy (harsimran_hs4):

well i think it satisfies 5x + y >= -10 5(0) + 0 = 0 0 >= -10 doesn`t it?

OpenStudy (anonymous):

Ooh .. yes it does so how would i shade it on the graph? @harsimran_hs4

OpenStudy (harsimran_hs4):

 the region A along with the line will satisfy the inequality

OpenStudy (anonymous):

Thank you so much!!! can you help me with one last question? @harsimran_hs4